OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −6(2u − 1)

Parth (parthkohli):

Have you tried it yet? :)

OpenStudy (anonymous):

yes im confused.

Parth (parthkohli):

Hmm. Where exactly? I'd love to help you!

OpenStudy (anonymous):

i dont understand how to do it with the variable in the parenthese... :)

OpenStudy (unklerhaukus):

\[a\times(b+c)=(a\times b)+(a\times c)\]

Parth (parthkohli):

Yes.

Parth (parthkohli):

\( \color{Black}{\Rightarrow -6(2u -1) = (-6)(2u)+ (- 6)(-1)}\)

OpenStudy (unklerhaukus):

\[−6(2u − 1)=(-6\times2u)+(-6\times-1)=\dots\]

OpenStudy (anonymous):

..... uhhh what?!

OpenStudy (unklerhaukus):

can you multiply the terms in the brackets, ?